A 0.410 L sample of argon is collected at 45 oC. If the temperature of the sample is increased to 75 oC, what volume will the gas occupy? Assume that the pressure and amount of gas are kept constant through the heating process.


1
Expert's answer
2021-07-21T04:39:59-0400

Given the initial volume of Argon gas(V1)=0.410L.

Initial Temperature(T1)="45\degree C"

=273+45=318K

T1=318K

Final Temperature (T2)="75\degree C"

=273+75=348K

T2=348K

Final Volume (V2)=?

Applying the Charles Law therefore,

"\frac{V_1}{T_1}=\frac{V_2}{T_2}"


"V_2=\frac{V_1×T_2}{T_1}"


V2="\frac{0.410L×348K}{318K}=0.450L"


Hence,the gas will occupy "0.450L"